# Global Hash Tables Strike Back! An Analysis of Parallel GROUP BY Aggregation

### v1

@@ 5. Related Work
The performance differences between concurrent aggregation using a shared hash table as opposed to a partitioned approach using a local aggregation table have been benchmarked and studied extensively over the years(Cieslewicz and Ross, 2007; Ye et al., 2011; Luan and Chang, 2022). None of these cited works propose leveraging indirection to reduce hash table operations to a lookup and insert-only workload, instead performing updates within the hash table itself, which requires significant concurrency control. As a result, the consensus in the literature is that thread local aggregation is not feasible due to contention costs in the presence of skew. We identified one work (Romanous et al., 2021) that also leverages a lookup and insert-only workload on a shared hash table by deferring update contention, but does so in the context of FPGAs and uses specialized hardware. Unlike this work, (Romanous et al., 2021) performs updates in the hash table but prevents simultaneous access using a hardware cache to synchronize all update requests on the same key. This cache is crucial to their approach and can not be adapted to general-purpose CPUs, but their findings align with ours that in providing a fast, contention-free path for lookups and preventing update contention in the hash table, far greater resistance to skew and high performance in low cardinality workloads can be achieved.
@@
Large amounts of research from outside the database domain on concurrent hash table designs are also relevant to our work. In Section 3.1 we tested state-of-the-art hash table designs from or related to those described in prior work (Li et al., 2014; Pandey et al., 2023; Maier et al., 2018; Kelly et al., 2020). However, as previously noted, even those papers that consider the use of concurrent hash tables for aggregation do not test them with regard to the specialized lookup and insert workload that underpins the ticketing step of our aggregation model, instead focusing on general-purpose workloads. While the body of work on concurrent hash tables is informative, each design must be critically reevaluated in the context of our specialized use case to make a determination about their performance for aggregations.
@@
In addition to the use of shared hash tables, many other methods for aggregation have been proposed. The method we most directly compare to is partitioning methods, which has been successfully integrated into many real-world systems (Leis et al., 2014; Kuiper et al., 2024; Lamb et al., 2023). Various hybrid approaches that leverage local aggregations for heavy hitters and a global shared table for other values have been proposed, in order to balance performance and memory concerns (Cieslewicz and Ross, 2007; Fent and Neumann, 2021). Ideas from these works may be beneficial for resolving the challenges we found with our fully concurrent aggregation at high thread count and cardinalities. Finally, sort-based aggregations have also been extensively evaluated (Vaghasiya and Jahangiri, 2024) but are less relevant to our work given that sort-based aggregation does not fit well into a morsel-driven model, as sorting is generally a pipeline breaker (although some variations like (Müller et al., 2015) propose methods that can act on smaller runs, but in a more limited capacity than intended).
@@
Finally, a very closely related operation to hash aggregations is hash joins, which has had a similar discourse about the merits of partitioning versus concurrent approaches. However, the nature of the hash join workload is significantly different, using hash tables as a multi-map rather than a map (e.g., the same key maps to multiple rows) to a partial aggregate. We leverage the fact that partial aggregates are densely stored in an underlying vector, an assumption that does not hold in hash aggregations. Furthermore, a large concern for hash joins is checking for the existence of a key within the probe-side relation, which is not relevant to the aggregation case. These unique performance characteristics have led to highly efficient concurrent hash table designs specialized for hash joins (Leis et al., 2014; Bandle et al., 2021; Birler et al., 2024). This body of work has had a similar overarching takeaway that specialized hash tables are crucial to achieving good performance in operators specific to the databases domain, but their methods are generally incompatible with the needs of fully concurrent aggregation.
@@ 6. Conclusions and future work
The go-to group by aggregation technique in modern analytic DBMSes is partitioning-based. In this work, we make the case for a simpler approach that takes advantage of a global concurrent hash table. In order for this simpler approach to match the performance of purpose-built partitioning techniques, a purpose-built hash table is required. We explored the operational benefits and drawbacks of both approaches and offered guidance to implementers along the way.
In future work, we plan to investigate more complex hash table designs as well as hybrid aggregation strategies (e.g., with teams of threads sharing an aggregation table to reduce contention). Another promising direction for investigation is understanding how new hardware features–like vectorized atomic operations–could be used to accelerate group aggregation even further. Finally, the fully concurrent approach we analyze here suffers from issues with resizing and spilling, which are ripe areas for future work.

### v2

@@ 5. Related Work
The performance differences between concurrent aggregation using a shared hash table versus a partitioned approach using a local aggregation table have been benchmarked and studied extensively over the years (Cieslewicz and Ross, 2007; Ye et al., 2011; Luan and Chang, 2022). Notably, prior work does not propose indirection to reduce hash table operations to a lookup and insert-only workload, instead performing updates within the hash table itself, which requires significant concurrency control. As a result, the consensus in the literature is that thread local aggregation is not feasible due to contention costs in the presence of skew. One exception is (Romanous et al., 2021), which used indirection, but in the context of FPGAs. Unlike our work, (Romanous et al., 2021) performs updates in the hash table but prevents simultaneous access using a hardware cache (not adaptable to general-purpose CPUs) to synchronize all update requests on the same key.
@@
Research from outside the database domain on concurrent hash table designs are also relevant to our work. In Section 3.1, we tested state-of-the-art hash table designs from or related to those described in prior work (Li et al., 2014; Pandey et al., 2023; Maier et al., 2018; Kelly et al., 2020). However, as previously noted, even prior work that consider hash aggregation did not customize their tables to only support the required operations, instead focusing on general-purpose workloads. While the body of work on concurrent hash tables is informative, each design must be critically reevaluated in the context of our specialized use case to make a determination about their performance for aggregations.
@@
In addition to the use of shared hash tables, many other methods for aggregation have been proposed. The method we compare most directly to is partitioned methods, which has been successfully integrated into many real-world systems (Raman et al., 2013; Leis et al., 2014; Kuiper et al., 2024; Lamb et al., 2023). Various hybrid approaches that leverage local aggregations for heavy hitters and a global shared table for other values have been proposed in order to balance performance and memory concerns (Cieslewicz and Ross, 2007; Fent and Neumann, 2021). Ideas from these works may be beneficial for resolving the challenges we found with our fully concurrent aggregation at high thread counts and cardinalities. Sort-based aggregations, which are exceptionally efficient when data is already sorted (Lamb et al., 2012), have also been extensively evaluated (Vaghasiya and Jahangiri, 2024b, a; Graefe et al., 1994). Müller et al. (Müller et al., 2015) note that when aggregating keys larger than CPU registers, sorting and hashing are computationally similar. In the Google F1 system, where distinct value counts are especially difficult to estimate, Do et al. (Do et al., 2022) showed that specialized sorting techniques could be more robust than hash-based aggregation. Sort-based aggregations also have operational advantages in out-of-core databases, as an external sort is generally simpler than external hashing (Vaghasiya and Jahangiri, 2024b; Do et al., 2022).
While this work focused on morsel-driven systems, other execution models, such as operator-level parallelism (Graefe and McKenna, 1993; Lamb et al., 2023) and pipeline parallelism (Zaharia et al., 2010), are also popular. It seems plausible that different aggregation methods may be optimal under different execution models, and we leave such investigations to future work.
@@
Finally, a closely-related operation to hash aggregations is hash joins. Similar debates about partitioning have occurred in the literature (Bandle et al., 2021), and specialized hash tables (multi-maps) tend to have large performance benefits (Leis et al., 2014; Birler et al., 2024). This body of work has had a similar overarching takeaway that specialized hash tables are crucial to achieving good performance, but their methods are generally incompatible with the needs of fully concurrent aggregation (e.g., multi-maps, probing).
@@ 6. Conclusions and future work
The go-to group by aggregation technique in modern analytic DBMSes is partitioning-based. In this work, we make the case for a simpler approach using a global concurrent hash table. For this simpler approach to match the performance of purpose-built partitioning techniques, a purpose-built hash table is required. We explored the operational benefits and drawbacks of both approaches and offered guidance to implementers along the way.
In future work, we plan to investigate more complex hash table designs as well as hybrid aggregation strategies (e.g., with teams of threads sharing an aggregation table to reduce contention). Another promising direction for investigation is understanding how new hardware features–like vectorized atomic operations–could be used to accelerate group aggregation even further. Additionally, the fully concurrent approach we analyze here suffers from issues with resizing and spilling, which are ripe areas for future work. Finally, this work only examined hash aggregation in a morsel-driven execution model: investigations into sort-based techniques, along with other execution models, could also yield interesting results.